Determine the osmotic pressure of a solution prepared by dissolving 25 mg of K2SO4 in 2 litre of water at 25° C, assuming that it is completely dissociated.

Grade:11

2 Answers

Pankaj
9 years ago

Mass of K2SO4, w = 25 mg = 0.025g (use 1 g = 1000 mg )
Volume V = 2 liter
T = 25 + 273 = 298 K
The reaction of dissociation of K2SO4
K2SO4 → 2K+ + SO42-
Number if ions produced = 2+ 1 = 3
So van’t Hoff factor i = 3
Now
\pi = i\frac{w}{M}\frac{1}{v}RT
Substituting the values and calculationg we get,
π= 5.27 × 10-3 atm
ankit singh
3 years ago

Given, that mass of K2SO4, w=25 mg=0.025 g

Volume V=2 L
T=25+273 =298 K
The reaction of dissociation of K2SO4 is,
K2SO42K++SO42
 
Number of ions produced=3=van't Hoff factor
 
Hence, π=iCRT
π=3×1740.025×0.5×0.0821×298=5.27×103 atm